# A Survey on Large Language Model-Based Game Agents

### v3

@@ Abstract.
Game environments provide rich, controllable settings that stimulate many aspects of real-world complexity. As such, game agents offer a valuable testbed for exploring capabilities relevant to Artificial General Intelligence (Yannakakis2018AIGames, ). Recently, the emergence of Large Language Models (LLMs) provides new opportunities to endow these agents with generalizable reasoning, memory, and adaptability in complex game environments. This survey offers an up-to-date review of LLM-based game agents (LLMGAs) through a unified reference architecture. At the single-agent level, we synthesize existing studies around three core components: memory, reasoning, and perception–action interfaces, which jointly characterize how language enables agents to perceive, think, and act. At the multi-agent level, we outline how communication protocols and organizational models support coordination, role differentiation, and large-scale social behaviors. To contextualize these designs, we introduce a challenge-centered taxonomy linking six major game genres to their dominant agent requirements, from low-latency control in action games to open-ended goal formation in sandbox worlds. A curated list of related papers is available at: https://github.com/git-disl/awesome-LLM-game-agent-papers.
@@ 1. Introduction
Existing surveys (wang2023survey, ; gao2023large, ) touch on the topic from different angles yet largely treat games as a downstream application alongside dialogue, tool use, or web automation, leaving the field of LLM-based game agents (LLMGAs) underexplored. The complexity and openness of game environments distinguish them from narrowly defined tasks. For instance, while a web-based agent may complete a query or transaction through a handful of API calls, a sandbox game enables researchers to cultivate entire agent societies and allows agents to freely explore, interact, and build within physics-driven worlds. These environments afford a level of freedom that enables emergent behaviors far beyond constrained, task-oriented interactions. On the other hand, game-focused surveys (gallotta2024large, ; sweetser2024large, ) emphasize areas such as game development, educational applications, or content generation. They neither examine the design challenges specific to LLMGAs nor explore the role of games as environments for advancing interactive intelligence. As a result, a dedicated survey of LLMGAs as a distinct research area is still missing.
Scope and Contributions. To fill this gap, we present this survey with two primary objectives. First, we categorize existing LLMGA studies under a unified reference architecture, which integrates two complementary parts: an LLMGA framework that enables component-level analysis of a single agent, and a multi-LLMGA framework that captures communication and organization within populations of agents. The LLMGA framework abstracts common choices into three modules: memory system, reasoning mechanism, and perception–action interface, each associated with a fundamental challenge. For instance, within the memory system, working memory faces limitations of capacity and temporal consistency, while long-term memory centers on when and what observations to consolidate and how to structure them for effective retrieval. The multi-LLMGA framework provides a complementary perspective that examines how agents communicate, coordinate, and self-organize under constraints such as partial observability, limited bandwidth, and evolving social dynamics. It distinguishes between agent-level communication, which governs message generation, interpretation, and belief alignment, and organization-level structure, which shapes topology, role assignment, and collective stability.
@@
This survey focuses exclusively on LLM-based agents in game environments. We include papers that (i) employ an LLM or MLLM as a central decision-making component and (ii) involve interaction with a game or game-like environment. We exclude both traditional non-LLM game AI approaches (e.g., deep reinforcement learning, symbolic systems) and LLM agents applied in non-game domains such as dialogue, web navigation, or API tool use. To construct the paper corpus, we searched four sources: ACM Digital Library, IEEE Xplore, Google Scholar, and arXiv, for the period 1 Jan 2018 to 31 Jul 2025 using the Boolean string: (large language model OR LLM) AND (game OR environment). Duplicates and irrelevant studies were removed, and additional works were identified through citation snowballing. Given the rapid pace of this field, we also maintain a curated and continuously updated collection of relevant literature at https://github.com/git-disl/awesome-LLM-game-agent-papers.
@@ 2.1. LLM-based Game Agent (LLMGA) Framework
Cognitive science views intelligence as an integrated system in which perception–action, memory, and reasoning processes interact to produce adaptive behavior (newell1994unified, ; kotseruba202040, ). In line with this view, we find that existing studies on LLMGAs primarily introduce techniques that fall into three components: memory, reasoning, and perception–action (park2023generative, ; yao2022react, ; hu2024pokellmon, ). Building on this perspective, we categorize existing LLMGA studies under a unified framework that instantiates these cognitive principles through the three components. Figure 1(a) illustrates the overall architecture: a central LLM connects the three components in continuous interaction with the game environment. At each step of gameplay, the environment evolves and produces new observations, which the agent perceives, interprets, and acts upon, completing a closed perception–action loop.
@@ 2.3. Game Taxonomy for LLMGA Design
Clarke et al. (clarke2017video, ) critically examine how conventional video game genre classifications often mix orthogonal dimensions such as mechanics and player structures, thereby lacking conceptual clarity. Building on this insight, we ground our taxonomy in established game studies literature through a gameplay-oriented perspective, drawing on the top-level groupings from SteamDB (steamdb2025, ) and the classification proposed by Lee et al. (lee2014facet, ). To maintain coherence with existing LLM-based game agent studies, we merge narrower categories (e.g., driving/racing, fighting) and additionally include sandbox games, resulting in six major genres as depicted in Figure 1. Building on this categorization, we further introduce a challenge-centered view, where each genre is linked to the core design challenge that most strongly drives agent development.
@@
As shown in Table 1, we identify six representative game genres, each posing distinct design challenges for LLM-based agents. (1) Action games unfold in real time and emphasize reflexive control, such as aiming, dodging, or chaining combos under tight temporal constraints. The core challenge is low-latency control, which shapes agent design by requiring fast action and hybrid architectures that reconcile LLM reasoning with frame-level responsiveness; (2) Adventure games emphasize exploration and long-horizon quests, where progress depends on remembering locations, items, and unresolved preconditions. The challenge is stateful world modeling, pushing agents to develop memory structures that maintain coherent records of evolving environments and dependencies; (3) Role-playing games center on character embodiment, where players assume predefined roles with distinct traits and narrative trajectories. The key challenge is role fidelity, shaping agent design toward embedding role profiles into memory and reasoning so that dialogue and actions remain persona-consistent over extended horizons; (4) Strategy games involve multi-step planning against adaptive adversaries, ranging from fully observable board games to imperfect-information settings with hidden states. Their central challenge is opponent-aware planning, which requires agents to integrate multi-step reasoning with theory-of-mind style opponent modeling; (5) Simulation games approximate real-world or systemic processes, from individual social life to the evolution of societies. The challenge is real-world fidelity, shaping agent design to ensure that behaviors remain credible and human-like rather than drifting into unrealistic patterns; (6) Sandbox games offer open-ended environments where players set their own objectives, explore, and build. The challenge is open-ended goal progression, which drives designs where agents can generate self-directed goals, decompose them hierarchically, and accumulate reusable skills to sustain long-term play.
@@ 3. Memory System of LLMGA
In cognitive psychology, working memory functions as a transient and limited-capacity buffer that temporarily stores and manipulates information needed for ongoing cognitive processing (BaddeleyHitch1974, ; baddeley2012working, ). In LLMGAs, this role is fulfilled by the model’s short context window and auxiliary mechanisms that keep recent observations “in mind” (shan2025cognitive, ). For working memory, we examine three key mechanisms. The first is context extension, which enlarges the effective context window so that recent events can be accommodated within short-term processing. The second is memory compression, which condenses lengthy inputs into compact representations, reducing capacity limits while preserving essential content. The third is active maintenance, which explicitly preserves recent bindings, plans, and intermediate states, preventing short-horizon drift and inconsistency caused by temporal decay.
@@
Figure 2. Overview of memory system of LLMGAs.
@@ 3.1. Working Memory
A complementary direction is summarization, which organizes long contexts into multi-level abstractions. For example, the chain-of-summarization approach incrementally condenses game-state trajectories by segmenting the temporal sequence into short windows and recursively summarizing them into higher-level representations. This hierarchical compression enables the model to retain the strategic context of long games while keeping the effective input size within the context window (starcraftii, ). Methods such as NUGGET cluster adjacent tokens into higher-level semantic “nuggets”, compact representations that compress contiguous text segments for efficient retrieval and long-context reasoning (qin2023nugget, ). WDMM constructs a memory tree and traverses it iteratively to surface only the most relevant segments (chen2023walking, ). These approaches echo the cognitive strategy of chunking, in which humans reduce information load by recoding multiple stimuli into higher-order units, thereby overcoming the intrinsic limits of working memory (cowan2001magical, ).
@@ 3.2. Long-Term Memory
Recent agent architectures emphasize three fundamental processes in the design of long-term memory systems. First, memory consolidation determines when and what to commit from transient buffers to durable storage, often triggered by event boundaries, importance scoring, or successful event execution (park2023generative, ; memorybank, ). Second, memory structuring addresses how stored content is organized, whether as raw chunks, key–value stores, hierarchical trees, knowledge graphs, or even implicit parametric memories fine-tuned into the model (wang2023voyager, ; anokhin2024arigraph, ). Finally, memory retrieval specifies how past knowledge is re-activated to guide ongoing decision-making, leveraging metadata filtering, semantic search, or traversal of graph/tree structures (li2024graphreader, ; park2023generative, ). These components together ensure that long-term memory efficiently and effectively archives past experience and supports future behavior.
@@
Finally, for parametrical storage, where knowledge is embedded implicitly in model weights rather than explicit structures. Such retrieval resembles implicit or procedural memory in humans, in which skills and habits are expressed without deliberate recall (shao2023character, ).
@@
LLMGA Enviornment Working Memory Long-Term Memory
@@ 4.1. Instruction-Guided Reasoning
Reflective Reasoning. Unlike generic LLM agents often evaluated on single-turn tasks, game agents operate within an observation–action–feedback loop, continuously perceiving the environment, taking actions, and adjusting decisions based on the resulting outcomes. Reflective reasoning builds on this loop by allowing agents to analyze the outcomes of their own actions, and incorporate these reflections into future reasoning and behavior, as Reflexion (shinn2023reflexion, ) shown in Figure 6. This introduces a temporal dimension to reasoning, enabling the integration of experience over time.
@@ 4.2. Fine-tuning for Improving Reasoning
In this subsection, we examine training paradigms for optimizing reasoning and action generation. Based on the training strategy, existing methods can be grouped into three categories. Supervised fine-tuning learns from expert trajectories to imitate observation–reasoning–action sequences. Reinforcement learning updates policies with reward feedback, reinforcing reasoning and actions that lead to favorable outcomes. Preference optimization leverages comparisons between better and worse trajectories to align models without the need for explicit reward models.
It is worth noting that some methods optimize only the final action without explicit reasoning, resembling what cognitive science terms System 1 thinking, i.e., fast, outcome-driven responses. However, they can be extended to System 2 thinking by eliciting chain-of-thought during interaction, allowing reasoning to be shaped through its effect on action outcomes (Kahneman2011, ).
@@
Recent works further integrate explicit reasoning into RL training, where the LLM is trained as a policy $\pi(rs_{t},a_{t}\mid s_{t})$. Reinforced Fine-Tuning (ReFT) (trung2024reft, ) introduces chain-of-thought supervision into PPO, encouraging the model to generate reasoning paths that lead to correct answers. However, because reasoning tokens are often much longer than action tokens, naive optimization can overweight reasoning relative to actions. Zhai et al. (zhai2025fine, ) propose downscaling the likelihood of reasoning steps, showing that moderate scaling achieves better balance between planning and acting. Beyond policy optimization, value-based methods such as Q-learning (mnih2013playing, ) extend RL to reasoning by treating partial generations as states and token expansions as actions. This formulation allows the use of search algorithms, such as Best-of-N sampling, Monte Carlo tree search, or A* search, to evaluate and expand reasoning paths guided by the Q-function (Q*, ; Q*Agent, ; dvo, ).
@@ 5. Perception and Action Interfaces of LLMGA
LLMGAs differ from generic LLM systems in that they operate within a continuous perception–action loop. To support this loop, they rely on a perception and action interface that serves as the agent’s API for interacting with the environment. On the input side, perception determines how raw game states, whether textual, symbolic, or visual, are abstracted into representations that can be processed by the LLM. On the output side, action generation ensures that the model’s decisions are translated into admissible in-game operations, ranging from natural language commands to sequences of low-level controls or programmatic calls. Figure 7 outlines the structure of this section. For the perception interface, we describe how to handle textual, symbolic, and visual observations; for the action interface, we explain how the outputs of LLMs are grounded into high-level, low-level, and code-based actions.
Figure 7. Summary of the perception and action interfaces.
@@ 5.2. Action Interface
Low-Level Actions. Low-level actions operate at the control layer, such as keystrokes, mouse movements, joystick inputs and are executed at each timestep. A low-level controller (policy) is responsible for translating a high-level action from the LLM into a sequence of control signals. One approach is heuristic planning (multi_agent_coordination, ; liu2023llm, ; park2023generative, ): given an intent such as “chop a tree,” the system invokes a path planner (e.g., A) to locate the nearest tree and then issues the necessary movement and interaction commands (zhu2023ghost, ). Another approach is to learn a low-level controller (policy) that generates the required action sequences to realize the LLM’s high-level decisions. Such policies can be trained either through imitation learning from expert demonstrations or through reinforcement learning with environment feedback, often aided by goal-conditioned rewards or semantic similarity between goals and observed transitions (rlgpt, ).

### v4

@@ Abstract.
Game environments provide rich, controllable settings that simulate many aspects of real-world complexity. As such, game agents offer a valuable testbed for exploring capabilities relevant to Artificial General Intelligence (Yannakakis2018AIGames, ). Recently, the emergence of Large Language Models (LLMs) provides new opportunities to endow these agents with generalizable reasoning, memory, and adaptability in complex game environments. This survey offers an up-to-date review of LLM-based game agents (LLMGAs) through a unified reference architecture. At the single-agent level, we synthesize existing studies around three core components: memory, reasoning, and perception–action interfaces, which jointly characterize how language enables agents to perceive, think, and act. At the multi-agent level, we outline how communication protocols and organizational models support coordination, role differentiation, and large-scale social behaviors. To contextualize these designs, we introduce a challenge-centered taxonomy linking six major game genres to their dominant agent requirements, from low-latency control in action games to open-ended goal formation in sandbox worlds. A curated list of related papers is available at: https://github.com/git-disl/awesome-LLM-game-agent-papers.
@@ 1. Introduction
Existing surveys (wang2023survey, ; gao2023large, ) touch on the topic from different angles yet largely treat games as a downstream application alongside dialogue, tool use, or web automation, leaving the field of LLM-based game agents (LLMGAs) underexplored. The complexity and openness of game environments distinguish them from narrowly defined tasks. For instance, while a web-based agent may complete a query or transaction through a handful of API calls, a sandbox game enables researchers to cultivate entire agent societies and allows agents to freely explore, interact, and build within physics-driven worlds. These environments afford a degree of freedom that enables emergent behaviors far beyond constrained, task-oriented interactions. On the other hand, game-focused surveys (gallotta2024large, ; sweetser2024large, ) emphasize areas such as game development, educational applications, or content generation. They neither examine the design challenges specific to LLMGAs nor explore the role of games as environments for advancing interactive intelligence. As a result, a dedicated survey of LLMGAs as a distinct research area is still missing.
Scope and Contributions. To fill this gap, we present this survey with two main objectives. First, we categorize existing LLMGA studies under a unified reference architecture, which integrates two complementary parts: an LLMGA framework that enables component-level analysis of a single agent, and a multi-LLMGA framework that captures communication and organization within populations of agents. The LLMGA framework abstracts common choices into three modules: memory system, reasoning mechanism, and perception–action interface, each associated with a fundamental challenge. For instance, within the memory system, working memory faces limitations of capacity and temporal consistency, while long-term memory centers on when and what observations to consolidate and how to structure them for effective retrieval. The multi-LLMGA framework provides a complementary perspective that examines how agents communicate, coordinate, and self-organize under constraints such as partial observability, limited bandwidth, and evolving social dynamics. It distinguishes between agent-level communication, which governs message generation, interpretation, and belief alignment, and organization-level structure, which shapes topology, role assignment, and collective stability.
@@
This survey focuses exclusively on LLM-based agents in game environments. We include papers that (i) employ an LLM or MLLM as a central decision-making component and (ii) involve interaction with a game or game-like environment. We exclude both traditional non-LLM game AI approaches (e.g., deep reinforcement learning, symbolic systems) and LLM agents applied in non-game domains such as dialogue, web navigation, or API tool use. To construct the paper corpus, we searched four sources: ACM Digital Library, IEEE Xplore, Google Scholar, and arXiv—for the period 1 Jan 2018 to 31 Jul 2025 using the Boolean string: ("large language model" OR LLM) AND (game OR environment). Duplicates and irrelevant studies were removed, and additional works were identified through citation snowballing. Given the rapid pace of this field, we also maintain a curated and continuously updated collection of relevant literature at https://github.com/git-disl/awesome-LLM-game-agent-papers.
@@ 2.1. LLM-based Game Agent (LLMGA) Framework
Cognitive science views intelligence as an integrated system in which perception–action, memory, and reasoning processes interact to produce adaptive behavior (newell1994unified, ; kotseruba202040, ). In line with this view, we find that existing studies on LLMGAs primarily introduce techniques that fall into three components: memory, reasoning, and perception–action (park2023generative, ; yao2022react, ; hu2024pokellmon, ). Building on this perspective, we categorize existing LLMGA studies under a unified framework that instantiates these cognitive principles through the three components. Figure 1(a) illustrates the overall architecture: a central LLM connects the three components in continuous interaction with the game environment. At each step of gameplay, the environment evolves and produces new observations, which the agent perceives, interprets, and acts upon, completing a closed perception-action loop.
@@ 2.3. Game Taxonomy for LLMGA Design
Clarke et al. (clarke2017video, ) critically examine how conventional video game genre classifications often mix orthogonal dimensions such as mechanics and player structures, thereby lacking conceptual clarity. Building on this insight, we ground our taxonomy in established game studies literature through a gameplay-oriented perspective, drawing on the top-level groupings from SteamDB (steamdb2025, ) and the classification proposed by Lee et al. (lee2014facet, ). To maintain coherence with existing LLMGA studies, we merge narrower categories (e.g., driving/racing, fighting) and additionally include sandbox games, resulting in six major genres as depicted in Figure 1. Building on this categorization, we further introduce a challenge-centered view, where each genre is linked to the core design challenge that most strongly drives agent development.
@@
As shown in Table 1, we identify six representative game genres, each posing distinct design challenges for LLM-based agents. (1) Action games (bellemare2013arcade, ; llm-colosseum, ) unfold in real time and emphasize reflexive control, such as aiming, dodging, or chaining combos under tight temporal constraints. The core challenge is low-latency control, which shapes agent design by requiring fast action and hybrid architectures that reconcile LLM reasoning with frame-level responsiveness; (2) Adventure games (wang2022scienceworld, ; hausknecht2020interactive, ) emphasize exploration and long-horizon quests, where progress depends on remembering locations, items, and unresolved preconditions. The challenge is stateful world modeling, pushing agents to develop memory structures that maintain coherent records of evolving environments and dependencies; (3) Role-playing games (wereworlf1, ; diplomacy, ) center on character embodiment, where players assume predefined roles with distinct traits and narrative trajectories. The key challenge is role fidelity, shaping agent design toward embedding role profiles into memory and reasoning so that dialogue and actions remain persona-consistent over extended horizons; (4) Strategy games (hupokellmon, ; starcraftii, ) involve multi-step planning against adaptive adversaries, ranging from fully observable board games to imperfect-information settings with hidden states. Their central challenge is opponent-aware planning, which requires agents to integrate multi-step reasoning with theory-of-mind style opponent modeling; (5) Simulation games (park2023generative, ; lin2023agentsims, ) approximate real-world or systemic processes, from individual social life to the evolution of societies. The challenge is real-world fidelity, shaping agent design to ensure that behaviors remain credible and human-like rather than drifting into unrealistic patterns; (6) Sandbox games (minecraft, ; hafner2021benchmarking, ) offer open-ended environments where players set their own objectives, explore, and build. The challenge is open-ended goal progression, which drives designs where agents can generate self-directed goals, decompose them hierarchically, and accumulate reusable skills to sustain long-term play.
@@ 3. Memory System of LLMGA
In cognitive psychology, working memory functions as a transient and limited-capacity buffer that temporarily stores and manipulates information needed for ongoing cognitive processing (BaddeleyHitch1974, ; baddeley2012working, ). In LLMGAs, this role is fulfilled by the model’s short context window and auxiliary mechanisms that keep recent observations “in mind” (shan2025cognitive, ). For working memory, we examine three key mechanisms. The first is context extension, which enlarges the effective context window so that recent events can be accommodated within short-term processing. The second is memory compression, which condenses lengthy inputs into compact representations, reducing capacity limits while preserving essential content. The third is active maintenance, which explicitly preserves recent bindings, plans, and intermediate states, preventing short-term drift and inconsistency caused by temporal decay.
@@
Figure 2. Overview of the memory system of LLMGAs.
@@ 3.1. Working Memory
A complementary direction is summarization, which organizes long contexts into multi-level abstractions. For example, the chain-of-summarization approach (starcraftii, ) incrementally condenses game-state trajectories by segmenting the temporal sequence into short windows and recursively summarizing them into higher-level representations. This hierarchical compression enables the model to retain the strategic context of long games while keeping the effective input size within the context window. Methods such as NUGGET cluster adjacent tokens into higher-level semantic “nuggets”, compact representations that compress contiguous text segments for efficient retrieval and long-context reasoning (qin2023nugget, ). WDMM constructs a memory tree and traverses it iteratively to surface only the most relevant segments (chen2023walking, ). These approaches echo the cognitive strategy of chunking, in which humans reduce information load by recoding multiple stimuli into higher-order units, thereby overcoming the intrinsic limits of working memory (cowan2001magical, ).
@@ 3.2. Long-Term Memory
Recent agent architectures emphasize three fundamental processes in the design of long-term memory systems. First, memory consolidation determines when and what to commit from transient buffers to durable storage, often triggered by event boundaries, importance scoring, or successful event execution (park2023generative, ; memorybank, ). Second, memory structuring addresses how stored content is organized, whether as raw chunks, key–value stores, hierarchical trees, knowledge graphs, or even implicit parametric memories fine-tuned into the model (wang2023voyager, ; anokhin2024arigraph, ). Finally, memory retrieval specifies how past knowledge is re-activated to guide ongoing decision-making, leveraging metadata filtering, semantic search, or traversal of graph/tree structures (li2024graphreader, ; park2023generative, ). These components together ensure that long-term memory effectively archives past experience and supports future behavior.
@@
Finally, for parametric storage, knowledge is embedded implicitly in model weights rather than explicit structures. Such retrieval resembles implicit or procedural memory in humans, in which skills and habits are expressed without deliberate recall (shao2023character, ).
@@
LLMGA Environment Working Memory Long-Term Memory
@@ 4.1. Instruction-Guided Reasoning
Reflective Reasoning. Unlike generic LLM agents often evaluated on single-turn tasks, game agents operate within an observation–action–feedback loop, continuously perceiving the environment, taking actions, and adjusting decisions based on the resulting outcomes. Reflective reasoning builds on this loop by allowing agents to analyze the outcomes of their own actions and incorporate these reflections into future reasoning and behavior, as Reflexion (shinn2023reflexion, ) shown in Figure 6. This introduces a temporal dimension to reasoning, enabling the integration of experience over time.
@@ 4.2. Fine-tuning for Improving Reasoning
In this subsection, we examine fine-tuning techniques for optimizing reasoning and action generation. Based on the training strategy, existing methods can be grouped into three categories. Supervised fine-tuning learns from expert trajectories to imitate reasoning and action generation. Reinforcement learning updates policies with reward feedback, reinforcing reasoning and actions that lead to favorable outcomes. Preference optimization leverages comparisons between better and worse trajectories to align models without the need for explicit reward models. It is worth noting that some methods mentioned below optimize only the final action without explicit reasoning, however, they can be extended to improve reasoning by eliciting chain-of-thought, allowing reasoning to be shaped through its effect on action outcomes (Kahneman2011, ).
@@
Recent works further integrate explicit reasoning into RL training, where the LLM is trained as a policy $\pi(rs_{t},a_{t}\mid s_{t})$. Reinforced Fine-Tuning (ReFT) (trung2024reft, ) introduces chain-of-thought supervision into PPO, encouraging the model to generate reasoning paths that lead to correct answers. However, because reasoning tokens are often much longer than action tokens, naive optimization can overweight reasoning relative to actions. Zhai et al. (zhai2025fine, ) propose downscaling the likelihood of reasoning steps, showing that moderate scaling achieves better balance between planning and acting. Beyond policy optimization, value-based methods such as Q-learning extend RL to reasoning by treating partial generations as states and token expansions as actions. This formulation allows the use of search algorithms, such as Best-of-N sampling or Monte Carlo tree search, to evaluate and expand reasoning paths guided by the Q-function (Q*, ; Q*Agent, ).
@@ 5. Perception and Action Interfaces of LLMGA
LLMGAs differ from generic LLM systems in that they operate within a continuous perception-action loop. To support this loop, agents rely on perception and action interfaces that serve as their eyes and hands for interacting with the environment (wang2023voyager, ; hu2024pokellmon, ). On the input side, the perception interface determines how raw game states are abstracted into representations that can be processed by the LLM, handling textual, symbolic, and visual observations. On the output side, the action interface ensures that the model’s decisions are translated into admissible in-game operations by grounding the LLM outputs into high-level, low-level, and code-based actions. Figure 7 outlines the structure of this section.
Figure 7. Overview of perception and action interfaces in LLMGAs.
@@ 5.2. Action Interface
Low-Level Actions. Low-level actions operate at the control layer, such as keystrokes, mouse movements, joystick inputs, and are executed at each timestep. A low-level controller (policy) is responsible for translating a high-level action from the LLM into a sequence of control signals. One approach is heuristic planning (multi_agent_coordination, ; liu2023llm, ; park2023generative, ): given an intent such as “chop a tree,” the system invokes a path planner (e.g., A) to locate the nearest tree and then issues the necessary movement and interaction commands (zhu2023ghost, ). Another approach is to learn a low-level controller (policy) that generates the required action sequences to realize the LLM’s high-level decisions. Such policies can be trained either through imitation learning from expert demonstrations or through reinforcement learning with environment feedback, often aided by goal-conditioned rewards or semantic similarity between goals and observed transitions (rlgpt, ).